OpenStudy (anonymous):

How do I set the problem a/b+c if a/b=5;b/c=2

OpenStudy (anonymous):

is this \[\frac{a}{b+c}\] we're trying to find?

OpenStudy (anonymous):

from a/b=5 we have a=5b from b/c=2 we have c=b/2 now for a/(b+c) we have 5b/(b+b/2)=10/3.. is that that the answer? i assumed it a/(b+c) rather than a/b + c

OpenStudy (anonymous):

that is correct if that's what we are trying to find

OpenStudy (anonymous):

i don't think we would be able to find out a/b + c from these conditions.. because for that we require value of c separately which we can't find out because we have three unknowns but only two conditions provided.. so i agree with @eigenschmeigen we have no option other than taking it as a/(b+c)...

OpenStudy (anonymous):

That's somewhat, what I got 10/7, could you brake it down so I can see where I went wrong. This a question in a teacher certification test.

OpenStudy (anonymous):

@thx1138 first confirm us if it was a/(b+c) to find out so that we can proceed..

OpenStudy (anonymous):

any way we have put the value of a as 5b in numerator and c as b/2 in denominator.. can u show us your steps..?

OpenStudy (anonymous):

Its a over b+c

OpenStudy (anonymous):

Wish I could its a little hard figuring out this site on a tablet.

OpenStudy (anonymous):

a/(b+c)= 5b/(b+(b/2))= 5b/((2b+b)/2)= (5b*2)/3b=10/3